Two people are jogging in a circular track in the same direction. One person an go completely around the reach in 21 minutes. The second person takes 18 minutes. If they both start running in the same place at the same time, how long will it take them to be together at this place if they continue to run?
Answer by Edwin McCravy(20055)

We find the LCM of 18 and 21
Line up the prime factorizations,
then bring them down to the bottom line
and multiply: 

 18 = 2×3×3
 21 =     3×7
-------------
LCM = 2×3×3×7 = 126

Answer = 126 minutes [which is 6 complete laps for 
                      the slower runner, since 6×21=126
                      and 7 complete laps for the faster 
                      runner, since 7×18 also = 126.

Edwin